OpenStudy (anonymous):

2. Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation.

OpenStudy (anonymous):

wait up

OpenStudy (anonymous):

3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 )

OpenStudy (anonymous):

3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

@cj49

OpenStudy (cj49):

3(3x-6) = 4(7x-1) 9x-18 = 28x-4(instead of 1 it sud be 4)